OpenStudy (anonymous):

Given the point (5, 6) and the slope of 5, find y when x = 21. A)86 B)90 C)105 D)111

OpenStudy (anonymous):

i think it's C but i just want confirmation and if not then an explanation on how to solve it please.

OpenStudy (08surya):

y-y1=m(x-x1) put m=5 and x1=5,y1=6,x=21 and calculate for y

OpenStudy (anonymous):

ty